a boat which has a speed of 5km/h in still water crosses a river of width 1 km along the shortest possible path in 15 minutes.the velocity of the river was in km/h is??

Grade:12

1 Answers

Tanuja
7 years ago
The answer is 3kmph because Time is width / 5sinx15/60 =1/5sinxSinx= 4/5. Cosx=3/5 so Velocity of river is 5 cosx = 3